Divide and proof my answer 5 17/20÷ 2 1/4
Answer by checkley75(3666)

5 17/20=(5*20+17)/20=117/20
2 1/4=(2*4+1)/4=(8+1)/4=9/4
(117/20)/(9/4)=117/20*4/9 = 117*4/20*9 468/180 NOW REDUCE BY DIVIDING BOTH NUMERATOR & DENOMINATOR BY 4&9 OR 36 WE GET:
13/5